OpenStudy (anonymous):

i need the step by step instructions please 11304 base6 divided by4 base 6

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (amistre64):

base 6; are these logs?

OpenStudy (amistre64):

\(\frac{\log_6(11304)}{\log_6(4)}\) like this?

OpenStudy (amistre64):

if so its a change of base from: \[log_4(11304)\]

OpenStudy (anonymous):

if u convert base 6 to base 10 the answer is 406 base 10

OpenStudy (amistre64):

6.7323... ?

OpenStudy (amistre64):

oh i see.... it might be talking about number systems, like binary, octinary, geritolinary and those lol

OpenStudy (anonymous):

yes i think so

OpenStudy (amistre64):

it hard to tell with such scant information... but you could be right, and prolly are :)

OpenStudy (anonymous):

the correct answer is 1514 base6 but i need to know how to get there

OpenStudy (amistre64):

base 6: 46656 7776 1296 216 36 6 right?

OpenStudy (amistre64):

0 1 2 3 4 5 10 11 12 13 14 15 20 21 22 23 24 25 30 31 32 33 34 ..........

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if you convert 1514 base 6 to base 10 it gives 1514 but how do you do thereverse

OpenStudy (anonymous):

it gives 406 i mean

OpenStudy (amistre64):

46656 7776 1296 216 36 6 1 is the key . .... 10000 1000 100 10 1 somehow lol

OpenStudy (amistre64):

6rt(11304) ??

OpenStudy (amistre64):

its between 7776 and 1296

OpenStudy (anonymous):

yea

OpenStudy (amistre64):

1*6^4 1^6^3 3^6^2 0*6^1 4*6^0 ??

OpenStudy (amistre64):

(1 * (6^4)) + (1 * (6^3)) + (3 * (6^2)) + (0 * (6^1)) + (4 * (6^0)) = 1 624

OpenStudy (amistre64):

and 4(6^0) = 4 lol

OpenStudy (anonymous):

this gives you its equivalent in base 19 which is 1604 and / 4 gives 406

OpenStudy (amistre64):

1624/4 = 406

OpenStudy (anonymous):

yeah - thats the answer in base 10

OpenStudy (amistre64):

/6^0? the /6^1 .. maybe?

OpenStudy (anonymous):

but i cant remember how yo convert this to bas e 6

OpenStudy (amistre64):

1296 216 36 6 1 406 - 216

OpenStudy (anonymous):

its 11304 base6 / 4 base6= 1514 base6 but how is it done

OpenStudy (amistre64):

n0000

OpenStudy (anonymous):

good question!!

OpenStudy (amistre64):

/216?

OpenStudy (amistre64):

1 + Remainder/36 + R/6 ??

OpenStudy (amistre64):

406/6^3 + R/6^2 +R/6 +R maybe?? R being the remainder

OpenStudy (anonymous):

yes i think u r on to it

OpenStudy (amistre64):

406/6^3 = 1 R.190 190/6^2 = 5 R.10 10/6^1 = 1 R.4

OpenStudy (amistre64):

4/6^0 = 4

OpenStudy (anonymous):

yep - well done

OpenStudy (amistre64):

i knew it had to be something lol

OpenStudy (amistre64):

throw stuff at it till it sticks :)

OpenStudy (anonymous):

i did this stuff many years ago when i did 'modern maths'

OpenStudy (amistre64):

yeah, I prolly did this back in finite math.... 15 years ago lol

OpenStudy (anonymous):

those answers are wrong

OpenStudy (anonymous):

1514 is wrong?

OpenStudy (anonymous):

i suppose it was a roundabout way to do it - the comversions to base 10 and back to 6 but im sure that 1514 base 6 is correct

OpenStudy (anonymous):

multiply 1514 base6 times 4 base 6

OpenStudy (anonymous):

right - it comes to 11304 base 6

OpenStudy (anonymous):

okay how is it divided

OpenStudy (anonymous):

don't forget u r dealing with base 6 here so when u multippy 4 x 4 it comes to 14 and carry 1 etc6 base 10 but this is 16/6 R 4 so u write

OpenStudy (anonymous):

4 and carry for the next multiplication

OpenStudy (anonymous):

well we naturally multiply using base 10 because thats what our nuber system is based on so it is easier (for me anyway) to convert to base 10 , perform the division in base 10 thenfinally convert back to base 6

OpenStudy (anonymous):

to summarise: 11304 - convert to base 6 working from rt to lft 4 X 6^0 = 4 0 x 6^1 = 0 3 x 6^2 = 108 1 x 6^3 = 216 1 x 6^4 = 1296 adding these up gives 1624 dividing by 4 gives 406 now converting 406 back to base 6 comes to 1514 as amis demonstrated above

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ok